OpenStudy (anonymous):

true or falsee? y = 11^x is written as log11y = x.

OpenStudy (anonymous):

Definition of Log Log(x) = y b^(y) = x

OpenStudy (anonymous):

Take the log at base 11 of both sides of y = 11^(x) \[\log_{11} (y) = x\]

OpenStudy (anonymous):

The \[\log_{11} \] will cancel out on one side leaving x

OpenStudy (anonymous):

the other side will be log_11(y) so yes if it is log11(y) than it is correct